# Near-Optimal Approximation of Matrix Functions by Lanczos-FAN. Amsel, T. Chen, A. Greenbaum, C. Musco, and C. Musco

### v2

@@ Abstract
We study the widely used Lanczos method for approximating the action of a matrix function $f(\mathbf{A})$ on a vector $\mathbf{b}$ (Lanczos-FA). For the function $f(\mathbf{A})=\mathbf{A}^{-1}$, it is known that, when $\mathbf{A}$ is positive definite, the $\mathbf{A}$-norm error of Lanczos-FA after $k$ iterations matches the optimal approximation from the Krylov subspace of degree $k$ generated by $\mathbf{A}$ and $\mathbf{b}$. In this work, we ask whether Lanczos-FA also obtains similarly strong optimality guarantees for other functions $f$.
We partially answer this question by showing that, up to a multiplicative approximation factor, Lanczos-FA also matches the optimal approximation from the Krylov subspace for any rational function with real poles outside the interval containing $\mathbf{A}$’s eigenvalues. The approximation factor depends on the degree of $f$’s denominator and the condition number of $\mathbf{A}$, but not on the number of iterations $k$. This result provides theoretical justification for the excellent performance of Lanczos-FA on functions that are well approximated by rationals. Additionally, using different techniques, we prove that Lanczos-FA achieves a weaker notion of optimality for the functions $\mathbf{A}^{1/2}$ and $\mathbf{A}^{-1/2}$. Experiments confirm that our new bounds more accurately predict the convergence of Lanczos-FA than existing bounds, in particular those based on uniform polynomial approximation.
keywords:
matrix function approximation, Lanczos, Krylov subspace method, optimal approximation, rational functions
{MSCcodes}
65F60, 65F50, 68Q25
@@ 1 Introduction
Given a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, the Lanczos method for matrix function approximation (Lanczos-FA) is perhaps the most widely used method for approximating the action of a matrix function $f(\mathbf{A})$ on a vector $\mathbf{b}\in\mathbb{R}^{d}$. Here, $f:\mathbb{R}\rightarrow\mathbb{R}$ is a scalar function and $f(\mathbf{A})$ is defined as $f(\mathbf{A}):=\sum_{i=1}^{d}f(\lambda_{i})\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}$, where $\mathbf{A}$ has eigendecomposition $\mathbf{A}=\sum_{i=1}^{d}\lambda_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}$. Like other Krylov subspace methods, Lanczos-FA iteratively constructs an approximation to $f(\mathbf{A})\mathbf{b}$ from the Krylov subspace
@@
In particular, the $k$-th Lanczos-FA iterate is defined as
$\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b}):=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b},$ (2)
where $\mathbf{Q}=[\mathbf{q}_{1},\ldots,\mathbf{q}_{k}]$ is an orthonormal basis for the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ such that $\operatorname{span}\{\mathbf{q}_{1},\ldots,\mathbf{q}_{j}\}=\mathcal{K}_{j}(\mathbf{A},\mathbf{b})$ for all $j\leq k$, and $\mathbf{T}:=\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}$. This algorithm can be implemented efficiently in $O(dk+k^{2}\log k)$ time, plus the time to compute $k$ matrix-vector products with $\mathbf{A}$; see for instance [MMS18, Appendix A].
Besides Lanczos-FA, a number of algorithms which construct approximations to $f(\mathbf{A})\mathbf{b}$ from the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ have been developed [${}^{+}$02, ${}^{+}$20, CGMM23]. While many of these algorithms come with stronger theoretical guarantees than currently known for Lanczos-FA, their practical performance often seems to be worse (and rarely substantially better) than Lanczos-FA. Thus, despite being one of the oldest and simplest iterative methods for computing approximations to $f(\mathbf{A})\mathbf{b}$, Lanczos-FA remains popular in practice. The main goal of this paper is to improve our theoretical understanding of why the method performs so well.
@@ 1.1 Optimality guarantees for Krylov subspace methods
We are motivated by the remarkable observation that, for a wide variety of problem instances, Lanczos-FA converges almost as quickly as the best approximation to $f(\mathbf{A})\mathbf{b}$ that could be returned by any Krylov subspace method run for the same number of iterations. In particular, all Krylov subspace methods output approximations that lie in the span of $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, that is, approximations of the form $p(\mathbf{A})\mathbf{b}$ for a polynomial $p$ of degree less than $k$. Given a problem instance $(f,\mathbf{A},\mathbf{b},k)$, the best possible approximation returned by a Krylov method is [^1] We choose to measure error in the Euclidean norm, and thus define optimality with respect to this norm. While other norms have been considered in prior work on matrix-function approximation [CGMM23], the Euclidean norm is the simplest and most natural starting point.
@@
While, in general, no efficient algorithm for computing $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$ is known, as shown in Fig. 1, the solution returned by Lanczos-FA often nearly matches the error of this best approximation. We are motivated to ask whether, at least for some class of problems, Lanczos-FA satisfies the following strong notion of approximate optimality:
Definition 1.1 (Near Instance Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near instance optimal with parameters $C,c$ if
@@
Above, $\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})$ denotes the output of the method obtained from the degree $k$ Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, i.e., the output after $k$ iterations.
Definition 1.1 is a near-optimality condition parameterized by two values, $C$ and $c$. Typically we think of $C$ as $\geq 1$ (which allows for the algorithm to return error slightly larger than the optimal approximation), and of $c$ as being $\leq 1$ (which allows us to compare against the best solution over a slightly smaller Krylov subspace). A near instance optimal guarantee is stronger when $C$ and $c$ are closer to $1$.
Notably, the right hand side in Definition 1.1 depends on the entire problem instance, $(f;\mathbf{A},\mathbf{b},k)$, which is why we call the guarantee “instance optimal”. The error of the Krylov optimal iterate for a given $f$ can also be upper bounded by a less precise quantity that is independent of $\mathbf{b}$ and only weakly dependent on $\mathbf{A}$. Specifically, if $\Lambda$ is the spectrum of $\mathbf{A}$ and $[\lambda_{\textup{min}},\lambda_{\textup{max}}]=\{\mathbf{x}^{\mathsf{T}}\mathbf{A}\mathbf{x}:\|\mathbf{x}\|_{2}=1\}$ is the field of values (FOV) of $\mathbf{A}$, we have:
$\|f(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ $\leq\|\mathbf{b}\|_{2}\min_{\deg(p)<k}\left(\max_{x\in\Lambda}|f(x)-p(x)|\right)$ (4)
$\leq\|\mathbf{b}\|_{2}\min_{\deg(p)<k}\left(\max_{x\in[\lambda_{\textup{min}},\lambda_{\textup{max}}]}|f(x)-p(x)|\right).$ (5)
@@
Figure 1: Lanczos-FA error $\|f(\mathbf{A})-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ at each iteration for several functions/spectra. “FOV Optimal” is the right hand side of Definition 1.3 and “Spectrum Optimal” is the right hand side of Definition 1.2, both with $C=c=1$. “Instance Optimal” is the right hand side of Definition 1.1 with $C=c=1$, which is a lower bound for all Krylov subspace methods, including Lanczos-FA. Lanczos-FA performs nearly instance optimally on a wide range of problems, far better than the bound of Fact 1 predicts. This is easily seen in the bottom plots, which show the ratio of the error of the Lanczos-FA iterate and the Krylov optimal iterate, $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$.
These upper bounds on the error of $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$ inspire natural alternative definitions of optimality that relax Definition 1.1.
Definition 1.2 (Near Spectrum Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near spectrum optimal with parameters $C,c$ if
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq C\|\mathbf{b}\|_{2}\min_{\deg(p)<ck}\left(\max_{x\in\Lambda}|f(x)-p(x)|\right).$
Definition 1.3 (Near FOV Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near FOV optimal with parameters $C,c$ if
Each type of optimality tightly relaxes the previous one. Definition 1.1 implies Definition 1.2 and, as we show in Subsection A.1, for any $f,k$ and $\mathbf{A}$, there exists a “worst-case” choice of $\mathbf{b}$ for which their bounds coincide; likewise, Definition 1.2 implies Definition 1.3, and for any $f$ and $k$ there exists a “worst-case” choice of $\mathbf{A}$ and $\mathbf{b}$ for which their bounds coincide. We also note that, with some additional assumptions on $\mathbf{b}$, spectrum optimality implies instance optimality.
Lemma 1.4.
A proof of Lemma 1.4 appears in Subsection A.2. The bound is useful e.g., when $\mathbf{b}$ is a random vector. For example, when $\mathbf{b}$ has independent and identically distributed Gaussian entries, $\frac{\|\mathbf{b}\|_{2}}{\min_{j}|\mathbf{u}_{j}^{\mathsf{T}}\mathbf{b}|}=O(d^{3/2})$ with high probability.
The three notions of near optimality given in Definition 1.1, Definition 1.2, and Definition 1.3, provide a clear and intuitive framework for understanding the convergence behavior of Krylov subspace algorithms like Lanczos-FA by comparing them to this optimal iterate. For example, one standard bound on the convergence of Lanczos-FA is the following well-known FOV optimality guarantee:
Fact 1.
For all problem instances $(f,\mathbf{A},\mathbf{b},k)$ where $\mathbf{A}$ is symmetric [^2] While we only study real symmetric matrices, the Crouzeix-Palencia Theorem can be used to generalize Fact 1 to arbitrary square matrices [CP17]. The domain of the maximization in Definition 1.3 should be changed from $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ to the field of values of $\mathbf{A}$, and the parameter $C$ should be changed from $2$ to $2(1+\sqrt{2})$., Lanczos-FA is nearly FOV optimal with parameters $C=2,c=1$.
Proof 1.5.
See [Saa92], [OSV12], or [MMS18, Theorem 4.1].
While powerful in its generality, as shown in Fig. 1, this bound is loose in many cases because it does not take into account properties of $\mathbf{A}$ such as isolated or clustered eigenvalues. Our goal in this paper is to improve this bound by proving spectrum and instance near optimality guarantees for various classes of functions $f$. We believe that studying Lanczos-FA through this lens is a promising way to further our understanding of its remarkable success in practice.
1.2 Existing near optimality guarantees for Lanczos-FA
With this goal in mind, we note that only a few instance optimality or spectrum optimality results are known for Lanczos-FA. Most directly, instance optimality trivially holds when $f$ is a polynomial of degree $<k$. In this case, it is well known that $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})$ exactly applies $f(\mathbf{A})\mathbf{b}$; i.e., $\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}=0$ [Saa92]. A more interesting special case is when $f(x)=1/x$ and $\mathbf{A}$ is positive definite. Then, the Lanczos-FA algorithm is mathematically equivalent to the well-known conjugate gradient algorithm for solving a system $\mathbf{A}\mathbf{x}=\mathbf{b}$. This implies that Lanczos-FA is the optimal approximation in the Krylov subspace with respect to the $\mathbf{A}$-norm; that is,
where $\kappa(\mathbf{A})$ is the condition number $\lambda_{\textup{max}}/\lambda_{\textup{min}}$ of $\mathbf{A}$. That is, Definition 1.1 is satisfied with $C=\sqrt{\kappa(\mathbf{A})}$ and $c=1$.
Note that the best polynomial approximation to $1/x$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ already converges at a geometric rate [JM05]:
That is, FOV optimality (Fact 1) suffices to prove the exponential convergence of the conjugate gradient method. However, it is well understood that conjugate gradient (and equivalently Lanczos-FA) often converge even faster, which is explained theoretically by the stronger instance and spectrum optimality guarantees that the methods satisfy[BK01, LS13, CLS22]. This illustrates the value of such guarantees over weaker bounds like Fact 1 for understanding the behavior of Lanczos-FA in general.
If $\mathbf{A}$ is not positive definite, $\mathbf{T}$ may have an eigenvalue at or near to zero and the error of the Lanczos-FA approximation to $\mathbf{A}^{-1}\mathbf{b}$ can be arbitrarily large. In fact, the same is true for any function $f$ which is much larger on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ than on $\Lambda$, the set of eigenvalues of $\mathbf{A}$. However, while the Lanczos-FA iterates may be bad at some iterations, the overall convergence of Lanczos-FA with $f(x)=1/x$ is actually good. In particular, the convergence of Lanczos-FA can be related to the MINRES algorithm [CG96], which produces an optimal approximation to $\mathbf{A}^{-1}\mathbf{b}$ with respect to the $\mathbf{A}^{2}$-norm. This allows certain optimality guarantees for MINRES to be transferred to Lanczos-FA, even on indefinite systems; see Subsection A.3 for details. In particular, for every $k$, there exists $k^{*}\leq k$ such that [^3] For indefinite matrices, $\kappa(\mathbf{A})$ denotes the ratio of the largest eigenvalue magnitude to the smallest eigenvalue magnitude.
$\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}\leq\kappa(\mathbf{A})\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (7)
While (7) does not quite fit the form of Definition 1.1 because of the $k^{*}$ on the left side, it is similar in spirit. Also note that the dependence on $k$ in the prefactor $\mathrm{e}\sqrt{k}+{1}/{\sqrt{k}}$ is not of great significance. Indeed, the convergence of the optimal polynomial approximation to $1/x$ on two intervals bounded away from zero is geometric.
Beyond the situations described above, the only other near optimality guarantee for Lanczos-FA that we are aware of is for the matrix exponential. In particular, [DGK98, (45)] shows that the Lanczos-FA iterate satisfies the guarantee
Again, this bound does not quite fit Definition 1.1 due to the maximization over $s$. The authors of [DGK98] state, “It is not known whether the maximum over $s\in[0,t]$ in (45) can be omitted and $s$ set equal to $t$ in the right-hand side.”

### v3

# Nearly Optimal Approximation of Matrix Functions by the Lanczos Method

@@ Abstract
Approximating the action of a matrix function $f(\mathbf{A})$ on a vector $\mathbf{b}$ is an increasingly important primitive in machine learning, data science, and statistics, with applications such as sampling high dimensional Gaussians, Gaussian process regression and Bayesian inference, principle component analysis, and approximating Hessian spectral densities. Over the past decade, a number of algorithms enjoying strong theoretical guarantees have been proposed for this task. Many of the most successful belong to a family of algorithms called Krylov subspace methods. Remarkably, a classic Krylov subspace method, called the Lanczos method for matrix functions (Lanczos-FA), frequently outperforms newer methods in practice. Our main result is a theoretical justification for this finding: we show that, for a natural class of rational functions, Lanczos-FA matches the error of the best possible Krylov subspace method up to a multiplicative approximation factor. The approximation factor depends on the degree of $f(x)$’s denominator and the condition number of $\mathbf{A}$, but not on the number of iterations $k$. Our result provides a strong justification for the excellent performance of Lanczos-FA, especially on functions that are well approximated by rationals, such as the matrix square root.
@@ 1 Introduction
Given a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ with eigendecomposition $\mathbf{A}=\sum_{i=1}^{d}\lambda_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}$, the matrix function $f(\mathbf{A})$ corresponding to a scalar function $f:\mathbb{R}\rightarrow\mathbb{R}$ is defined as
$f(\mathbf{A}):=\sum_{i=1}^{d}f(\lambda_{i})\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}.$ (1)
Matrix functions arise throughout machine learning, data science, and statistics. For instance, the matrix square root is used in sampling Gaussians, Bayesian modeling, and Gaussian processes [aune_eidsvik_pokern_13, aune_simpson_eidsvik_13, pleiss_jankowiak_eriksson_damle_gardner_20], general fractional matrix powers are used in Markov chain modeling and Rényi entropy estimation [waugh_abel_67, janson_92, jarrow_lando_turnbull_97, dong_gong_yu_li_23], the matrix logarithm is used for determinantal point processes, kernel learning, and approximating log-determinants for Gaussian process regression and Bayesian inference [dong_eriksson_nickisch_bindel_wilson_17, han_malioutov_avron_shin_17, gardner_pleiss_weinberger_bindel_wilson_18], the matrix sign is used in principal components regression and spectral density estimation [frostig_musco_musco_sidford_16, dong_eriksson_nickish_bindel_wilson_17, jin_sidford_19, papyan_19, ghorbani_krishnan_xiao_19, yao_gholami_keutzer_mahoney_20, chen_trogdon_ubaru_21, braverman_krishnan_musco_22, chen_trogdon_ubaru_22], and the matrix exponential is used in network science [aprahamian_higham_higham_15, wang_sun_musco_bao_21, massei_tudisco_24]. In many of these applications, we do not need to compute the matrix $f(\mathbf{A})$ itself, but rather its action on a vector $\mathbf{b}\in\mathbb{R}^{d}$; i.e., $f(\mathbf{A})\mathbf{b}$. This task can be performed much more efficiently than computing an eigendecomposition of $\mathbf{A}$ and forming $f(\mathbf{A})$ using Equation 1.
Perhaps the first general purpose method for approximating $f(\mathbf{A})\mathbf{b}$ is the Lanczos method for matrix function approximation (Lanczos-FA) [druskin_knizhnerman_89, gallopoulos_saad_92], which is the focus of the present paper. Over the past decade, a number of special purpose algorithms, designed for a single function or class of functions, have been developed [eshof_frommer_lippert_schilling_van_der_vorst_02, cundy_krieg_arnold_frommer_lippert_schilling_09, frommer_kahl_schweitzer_tsolakis_23, jin_sidford_19, chen_greenbaum_musco_musco_23]. These newer algorithms often satisfy strong theoretical guarantees better than the best-known bounds for Lanczos-FA. The present paper is motivated by the following remarkable observation:
Despite being arguably the simplest and most general algorithm for computing $f(\mathbf{A})\mathbf{b}$, Lanczos-FA frequently outperforms special purpose algorithms, sometimes by orders of magnitude, on common test problems.
For instance, in Figures 4, 5, and 8 we compare Lanczos-FA with specialized algorithms [jin_sidford_19, pleiss_jankowiak_eriksson_damle_gardner_20, chen_greenbaum_musco_musco_23] that satisfy the best-known theoretical guarantees for computing $f(\mathbf{A})\mathbf{b}$ for the particular functions they were designed for. In these experiments Lanczos-FA drastically outperforms these methods, despite the fact that it was not designed for any particular function. Because of its outstanding performance, Lanczos-FA is perhaps the most commonly used algorithm for computing $f(\mathbf{A})\mathbf{b}$ in practice. The main goal of this paper is to improve our theoretical understanding of why Lanczos-FA performs so well, in order to help close the theory-practice gap.
1.1 Krylov subspace methods
Lanczos-FA falls into a class of algorithms called Krylov Subspace Methods (KSMs). KSMs are among the most powerful and widely used algorithms for a broad range of computational tasks including solving linear systems, computing eigenvalues/vectors, and low-rank approximation [saad_03, halko_martinsson_tropp_11, liesen_strakos_13, tropp_webber_23]. Like other KSMs for computing $f(\mathbf{A})\mathbf{b}$, Lanczos-FA iteratively constructs an approximation from the Krylov subspace
@@ 1.1 Krylov subspace methods
The Lanczos algorithm [lanczos_50] produces an orthonormal basis $\mathbf{Q}=[\mathbf{q}_{1},\ldots,\mathbf{q}_{k}]$ for the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and a symmetric tridiagonal matrix $\mathbf{T}$ satisfying $\mathbf{T}=\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}$ [saad_03]. The Lanczos-FA algorithm uses this $\mathbf{Q}$ and $\mathbf{T}$ to approximate $f(\mathbf{A})\mathbf{b}$.
Definition 1.
The Lanczos-FA iterate for a problem instance $(f,\mathbf{A},\mathbf{b},k)$ is defined as
$\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b}):=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b},$
where $\mathbf{Q}$ and $\mathbf{T}$ are as above.
In our analysis we assume exact arithmetic. We do not discuss the implementation of Lanczos or Lanczos-FA since there are many resources on this topic; see for instance [musco_musco_sidford_18, carson_liesen_strakos_24, chen_24]. Fortunately, if the Lanczos algorithm is implemented properly, its finite-precision behavior closely follows its exact-arithmetic behavior on a nearby problem [greenbaum_89]. See Section 5 for further discussion.
@@ 1.2 Optimality guarantees for Krylov subspace methods
For a wide variety of problem instances, Lanczos-FA is observed to converge almost as quickly as the best approximation to $f(\mathbf{A})\mathbf{b}$ that could be returned by any KSM run for the same number of iterations. In particular, all KSMs output approximations that lie in the span of $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, that is, approximations of the form $p(\mathbf{A})\mathbf{b}$ for a polynomial $p$ of degree less than $k$. Given a problem instance $(f,\mathbf{A},\mathbf{b},k)$, the best possible approximation returned by a Krylov method is [^1] We choose to measure error in the Euclidean norm, and thus define optimality with respect to this norm. While other norms have been considered in prior work on matrix-function approximation [chen_greenbaum_musco_musco_23], the Euclidean norm a simple starting point.
@@
For general matrix functions, no efficient algorithm for computing $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$ is known, but as shown in Figure 1, the solution returned by Lanczos-FA often nearly matches the error of this best approximation. It is thus natural to ask if, at least for some class of problems, Lanczos-FA satisfies the following strong notion of approximate optimality:
Definition 2 (Near Instance Optimality).
For a problem instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is nearly instance optimal with parameters $C$ and $c$ if
@@
Above, $\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})$ denotes the output of an algorithm (e.g., Lanczos-FA) obtained from the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, i.e., the output after $k$ iterations.
@@
Figure 1: Lanczos-FA error $\|f(\mathbf{A})-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ at each iteration for several functions/spectra. “Instance Optimal” is the right hand side of Definition 2 with $C=c=1$, which is a lower bound for all KSMs, including Lanczos-FA. Lanczos-FA performs nearly instance optimally on a wide range of problems, far better than Fact 3 predicts. This is easily seen in the bottom plots, which show the ratio of the error of the Lanczos-FA iterate and the Krylov optimal iterate, $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$.
In Definition 2, $C\geq 1$ and $c\leq 1$ allow some slack in comparing to the Krylov optimal iterate. The right hand side depends on the entire problem instance, $(f;\mathbf{A},\mathbf{b})$, which is why we call the guarantee “instance optimal”.
1.3 Existing near-optimality analyses of Lanczos-FA
The best bound for Lanczos-FA applying to a broad class of functions is the following common bound for Lanczos-FA; see for instance [saad_92, orecchia_sachdeva_nisheeth_12, musco_musco_sidford_18, chen_greenbaum_musco_musco_22].
Fact 3.
For all problem instances $(f,\mathbf{A},\mathbf{b},k)$, Lanczos-FA satisfies
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq 2\|\mathbf{b}\|_{2}\min_{\deg(p)<k}\left(\max_{x\in[\lambda_{\textup{min}},\lambda_{\textup{max}}]}|f(x)-p(x)|\right).$
Fact 3 is in some sense an optimality guarantee; it compares the convergence of Lanczos-FA to the best possible uniform polynomial approximation to $f(x)$. However, it does not take into account properties of $\mathbf{A}$ such as isolated or clustered eigenvalues, and as seen in Figure 1, it typically only gives a loose upper bound on the performance of Lanczos-FA.
To date, near-instance-optimality guarantees for Lanczos-FA akin to those of Definition 2 are known only in a few special cases. The most well-known is when $f(x)=1/x$ and $\mathbf{A}$ is positive definite, in which case Lanczos-FA is mathematically equivalent to the celebrated Conjugate Gradient algorithm, and therefore exactly optimal in the $\mathbf{A}$-norm [liesen_strakos_13]. The instance-optimality guarantee for Lanczos-FA in this setting is used to prove well-known super-exponential convergence in certain settings [beckermann_kuijlaars_01, liesen_strakos_13, carson_liesen_strakos_24]. In contrast, a bound like Fact 3 only provides exponential convergence and is widely understood by the numerical linear algebra community to not accurately describe the actual behavior of the algorithm in most cases [greenbaum_97, liesen_strakos_13, carson_liesen_strakos_24]. The only other near-optimality guarantees for Lanczos-FA of which we are aware concern $\mathbf{A}^{-1}\mathbf{b}$ for nonsymmetric $\mathbf{A}$ [chen_meurant_24] and the matrix exponential $\exp(-t\mathbf{A})\mathbf{b}$ [druskin_greenbaum_knizhnerman_98]. In both cases, Lanczos-FA satisfies guarantees that are reminiscent of (although weaker than) Definition 2. Further discussion is given in Section B.1.
We also remark that there are a number of works which aim to relate the convergence of of Lanczos-FA for functions with certain integral representations to the convergence of conjugate gradient on linear systems [ilic_turner_simpson_09, frommer_guttel_schweitzer_14, frommer_schweitzer_15, frommer_simoncini_09, frommer_kahl_lippert_rittich_13, chen_greenbaum_musco_musco_22]. While these analyses provides spectrum dependent convergence guarantees, they are weaker than Definition 2 because it is not clear how the best possible KSM approximation to a given function relates to the convergence of conjugate gradient. These bounds were mostly developed for use as a posteriori stopping criteria rather than as a theoretical explication for the behavior of Lanczos-FA.